### v1

# Bridging and Modeling Correlations in Pairwise Data for Direct Preference Optimization

Abstract
Direct preference optimization (DPO), a widely adopted offline preference optimization algorithm, aims to align large language models (LLMs) with human-desired behaviors using pairwise preference data. However, the winning response and the losing response within pairwise data are generated isolatedly, leading to weak correlations between them as well as suboptimal alignment performance. To address this issue, we propose an effective framework named BMC, for bridging and modeling correlations in pairwise data. Firstly, we increase the consistency and informativeness of the pairwise preference signals by targeted modifications, synthesizing a pseudo winning response through improving the losing response based on the winning response. Secondly, we identify that DPO alone is insufficient to model these correlations and capture nuanced variations. Therefore, we propose learning token-level correlations by dynamically leveraging the policy model’s confidence during training. Comprehensive experiments on QA, math, and instruction-following tasks demonstrate the effectiveness of our approach, significantly surpassing competitive baselines, including DPO. Additionally, our in-depth quantitative analysis reveals the reasons behind our method’s superior performance over DPO and showcases its versatility to other DPO variants. We release our repository at https://github.com/YJiangcm/BMC.
1 Introduction
Preference optimization is essential for aligning large language models (LLMs) with human values and intentions, e.g., helpfulness, honesty, and harmlessness (Askell et al., 2021; Ouyang et al., 2022). Reinforcement learning from human feedback (RLHF) (Christiano et al., 2017; Bai et al., 2022a; Ouyang et al., 2022) is a widely adopted technique for achieving this alignment effectively. Despite the impressive results of RLHF, it faces significant optimization challenges due to its multi-stage process, which includes training a reward model and subsequently optimizing a policy model to maximize this reward (Casper et al., 2023).
An effective and straightforward alternative to RLHF is direct preference optimization (DPO) (Rafailov et al., 2024b). Unlike the traditional RLHF approach, DPO bypasses training a reward model and avoids using any reinforcement learning algorithms. Instead, it reparameterizes the reward function to directly learn a policy model from offline pairwise preference data, employing the Bradley-Terry ranking objective (Bradley & Terry, 1952). However, the generation of winning and losing responses within preference data often occurs in an isolated manner, whether through human annotation (Bai et al., 2022a) or automated techniques such as RLAIF (Bai et al., 2022b) and reject sampling (Liu et al., 2023; Pace et al., 2024). This isolation implies that winning and losing responses are produced without mutual visibility, resulting in a lack of strong correlation or relevance between them. Consequently, this detachment hinders the capacity to discern subtle yet critical differences between responses, potentially leading to suboptimal optimization and model alignment.
In this work, we introduce an innovative framework, termed BMC, to bridge and model correlations in pairwise data for direct preference optimization. During the Bridging Phase, we enhance correlations by increasing the consistency and informativeness of pairwise preference signals. By using the preferred response as a reference, we synthesize a pseudo preferred response through targeted modifications of the dispreferred response. This pseudo preferred response offers two key advantages: (1) it maintains high correlations with the dispreferred response while encapsulating all human-desired values in the preferred response; and (2) it serves as an interpolation between the winning and losing responses, easing the learning process. The nuanced differences between the pseudo-preferred and dispreferred responses are indeed what we expect the model to learn in the subsequent Modeling Phase. Nonetheless, we identify that DPO alone is insufficient to model these correlations and capture nuanced variations. From the perspective of the token-level Markov Decision Process (MDP) (Rafailov et al., 2024a), DPO aggregates rewards uniformly across all tokens, assuming equal contribution to sequence quality and neglecting token-specific importance. To address this, we adjust the emphasis on rewards of different tokens between the pseudo-preferred and dispreferred responses. Unlike previous methods (Guo et al., 2023; Cao et al., 2024; Chan et al., 2024; Chen et al., 2024a) that assign predefined values for fine-grained guidance, our adjustment is dynamically guided by the policy model’s confidence, i.e., the probability assigned to generated tokens during training. This ensures the model focuses on learning challenging distinctions while reinforcing known patterns, resulting in a more nuanced and robust policy.
We conduct extensive experiments across three downstream scenarios: question answering, mathematical reasoning, and instruction following, utilizing a total of 10 datasets. Our results demonstrate that our method consistently and significantly outperforms competitive offline optimization algorithms across various tasks. Notably, our approach surpasses vanilla DPO by up to 3.8 points on question-answering tasks and by up to 6.4 points on AlpacaEval 2. Furthermore, we use in-depth analyses to elucidate why our method outperforms DPO and show that our framework can be versatilely adapted to other DPO variants, confirming its potential for broad application.
2 Related Work
Preference optimization.
Preference optimization refers to aligning large language models with human preferences or specific desired outcomes. A well-established method for this is reinforcement learning from human feedback (RLHF) (Christiano et al., 2017; Bai et al., 2022a; Ouyang et al., 2022), which typically involves three stages: supervised fine-tuning, reward model training, and policy optimization. Although RLHF produces highly capable models, its training process is often complex and unstable (Santacroce et al., 2023), requiring the training of multiple LMs and sampling from the LM policy. To address these challenges, Direct Preference Optimization (DPO) (Rafailov et al., 2024b) introduces an alternative offline algorithm to optimize the regularized expected rewards without relying on RL. Following the introduction of DPO, several studies have sought to enhance it in various ways. For instance, IPO (Azar et al., 2024) aims to prevent DPO from overfitting to the preference dataset; ORPO (Hong et al., 2024) and SimPO (Meng et al., 2024) eliminate the need for a reference model; and R-DPO (Park et al., 2024) incorporates a regularization term to prevent exploitation based on length.
Preference data construction.
Constructing high-quality pairwise preference data is essential for preference optimization. Given the high cost of manually curating these datasets at scale, researchers have explored automated methods for producing preference data. One notable approach, RLAIF (Bai et al., 2022b) employs LLMs to label side-by-side response pairs, eliminating the need for human labeling. Alternatively, positive and negative responses can be generated by utilizing models of varying quality (Kim et al., 2023) or through specific prompting techniques (Yang et al., 2023). Recently, sampling-based methods such as Statistical Rejection Sampling (Liu et al., 2023) and West-of-N (Pace et al., 2024) have been introduced, generating preference pairs by selecting candidates sampled from the optimal policy. Nonetheless, these methods independently generate preferred and undesired responses without accounting for the correlations between them.
Token-level preference optimization.
The majority of preference optimization strategies typically utilize trajectory-wise (sequence-level) rewards, while LM training and generation both occur at the token level (Yang et al., 2024). To bridge this gap, FIGA (Guo et al., 2023) and DRLC (Cao et al., 2024) exploit external LLMs to pinpoint positive and negative token segments within responses, assigning fixed reward values (e.g., +1 for positive, -1 for negative) as guidance. Meanwhile, ABC (Chan et al., 2024) and RLMEC (Chen et al., 2024a) extract fine-grained credits from the reward model. Despite their contributions, these methods rely on predefined values for fine-grained guidance, failing to account for the dynamic learning process of the policy model.
3 Methodology
In this section, we present the proposed approach BMC by bridging and modeling correlations in pairwise data for direct preference optimization. As depicted in Figure 1, our BMC framework is structured around two pivotal stages: (1) the Bridging Phase, where we enhance the correlations between pairwise data by increasing the consistency and informativeness of pairwise preference signals through targeted modifications (§3.1); and (2) the Modeling Phase, where we dynamically model the correlations during the optimization process by leveraging the confidence of the policy model (§3.2), alleviating the insufficient token-level credit assignment of DPO.
3.1 Bridging Phase
In offline preference optimization, it is commonly assumed that we have access to a static pairwise preference dataset , where and denote the preferred and dispreferred completion, give the input prompt . To align an LLM, a basic principle is to ensure that the distribution of the model should not experience significant shifts during the alignment process (Bai et al., 2022a). Despite that is human preferred, it is likely to span a very different semantic distribution as . Additionally, since and are generated isolatedly, the correlation between and can be weak during pairwise preference optimization. This weak correlation poses a challenge, as the preferred outputs may not provide sufficiently informative gradients for adjusting the model’s parameters in relation to . Consequently, the optimization process may struggle to effectively leverage the provided pairwise preferences to fine-tune the model, potentially resulting in suboptimal alignment performance. To address this challenge, we enhance the alignment efficacy by improving the consistency and informativeness of the pairwise preference signals. Specifically, we utilize an off-the-shelf LLM to make targeted modification to by referring to :
where is the generated pseudo preferred completion, is the instruction prompt that requires to be modified only on dispreferred tokens compared to . In this way, we can effectively reduce the distribution shifts as well as obtain demonstrations more correlated to . Finally, we obtain a new dataset for subsequent training.
3.2 Modeling Phase
The difference between and can be easily obtained through dynamic programming algorithms like the Levenshtein Distance (Yujian & Bo, 2007), which highlights the core human expected and unexpected behaviors by decoupling from the inherent linguistic style and overall semantic distribution. To enhance the model’s ability to discern underlying patterns and deviations in human behavior, stopwords like “the” and “a” are filtered out, allowing for a more focused analysis of the substantive elements of the text. Despite these improvements, our findings indicate that DPO alone is insufficient for capturing the nuanced variances, suggesting the need for additional methods to fully model these correlations.
Limitations of DPO.
DPO (Rafailov et al., 2024b) stands as one of the leading methods for offline preference optimization. Rather than learning an explicit reward model, DPO reparameterizes the reward function using a closed-form expression with the optimal policy:
where and signify the policy model and reference model, respectively. is the partition function. By substituting the reparameterization into the Bradley-Terry model (Bradley & Terry, 1952) , DPO directly optimizes for the policy best satisfying the preferences with a simple classification objective:
DPO was originally derived as a bandit problem in which the whole response of the model is treated as a single arm. Alternatively, it can be deduced in the token-level MDP (Rafailov et al., 2024a):
where and denote the win trajectory and the lose trajectory, respectively. indicates the action (current generated token), and signifies the state (all tokens generated so far).
From Eq. 4, it can be seen that DPO assigns rewards to each token generation by , and simply add up the rewards of all tokens as the accumulated reward of the trajectory. This uniform aggregation assumes that each token contributes equally to the overall quality of the sequence, without considering the varying importance of each token (timestep). Therefore, nuanced differences between and that significantly influence the overall meaning or quality of the response might not be adequately emphasized (refer to Figure 6), leading to suboptimal performance.
Our solution.
To this end, we propose to emphasize the rewards of critical tokens, i.e., nuanced differences between and . The magnitude of the emphasis is determined dynamically by the policy model’s confidence, which refers to the probability assigned to the generated token during training.
For varied tokens in the preferred response , low policy confidence signifies areas of insufficient learning, indicating a need for greater emphasis to enhance the model’s ability to discern these nuances. Consequently, we modify the reward factor to be inversely proportional to the policy model’s confidence, as demonstrated in Eq. 8.
For varied tokens in the dispreferred response , we adjust the reward factor by leveraging the autoregressive property of the policy model. Through the training of the policy model, it tends to assign a low probability to given the input . By analyzing token-level probabilities, we observe that the probability of the initial token in undesired spans is low, whereas the probabilities of subsequent tokens are significantly higher, as illustrated in Figure 2. This phenomenon arises from the autoregressive nature of the model, where subsequent mistakes are conditioned on previous errors. Although the onset of the undesired span should warrant penalization, the subsequent tokens might be contextually appropriate if the policy model’s confidence is high, justifying reduced penalization. Therefore, we modify the reward factor to be inversely proportional to the policy model’s confidence, as described in Eq. 11.
In a nutshell, the magnitude of the emphasis on critical tokens is determined dynamically by the policy model’s confidence. This adaptive mechanism ensures that the model focuses on learning challenging distinctions while reinforcing already learned patterns, leading to a more nuanced and robust policy. The formalization of this approach is encapsulated in Eq. 5, where and dynamically adjust based on the policy’s confidence, ensuring a tailored emphasis on critical tokens to improve the overall model performance.
where
The denotes the stop-gradient operator, the is an upper limit threshold that controls the emphasis on the rewards of the critical tokens, preventing overly aggressive updates. The and signify using the Levenshtein Distance algorithm to find the varied tokens in and , respectively.
Gradient analysis.
For a mechanistic understanding of our method, we analyze the gradient of the loss function . The gradient with respect to the parameters can be written as:
where . It can be seen that our method takes into account both the sequential level and the token level perspectives: (1) at the sequential level, we encourage the preferred completion while punishing the dispreferred one; and (2) at the token level, we further refine the rewards of critical desired and undesired tokens of and , respectively. This dual consideration ensures that both the overall sequence structure and the individual token choices are optimized for the desired outcome.
4 Experimental Setup
We conduct three downstream scenarios for a comprehensive evaluation, including question answering (QA), mathematical reasoning, and instruction following (IF). The detailed data statistics as well as the evaluation metrics are listed in Table 5 of Appendix A.1.
Models and training settings.
For the QA and mathematical reasoning setup, we utilize Llama2-7B-base (Touvron et al., 2023) in our experiments. Successfully solving these tasks necessitates LLMs to possess domain-specific knowledge and engage in systematic, step-by-step reasoning to reach the ultimate answer. Therefore, following prior works (Chen et al., 2024a; b), we fine-tune Llama2-7B-base on the training set of ECQA (Aggarwal et al., 2021) and QASC (Khot et al., 2020) for QA, and fine-tune Llama2-7B-base on MetaMathQA (Yu et al., 2023) for mathematical reasoning. We denote the fine-tuned LLM as SFT and use it as the backbone for preference optimization. In line with prior research (Chen et al., 2024a; b), we construct preference pairs based on the training data, by using the ground truth as and the SFT model’s inference output as .
For the instruction following setup, we utilize Llama3-8B-base (AI@Meta, 2024) and Mistral-7B-Base (Jiang et al., 2023) in our experiments. Following the training pipeline of Zephyr (Tunstall et al., 2023) and SimPO (Meng et al., 2024), we train a base model on the UltraChat-200k dataset (Ding et al., 2023) to obtain an SFT model. Then, we use the SFT model as the starting point and perform preference optimization on the UltraFeedback dataset (Cui et al., 2023), where and are collected from LLMs of varying quality.
Evaluation benchmarks.
In question answering, we adopt the test splits of ECQA (Aggarwal et al., 2021), QASC (Khot et al., 2020), OpenbookQA (Mihaylov et al., 2018), and StrategyQA (Geva et al., 2021) for evaluation. In mathematical reasoning, we conduct the evaluation on four challenge datasets including GSM8k (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), MAWPS (Koncel-Kedziorski et al., 2016), and TabMWP (Lu et al., 2023). In instruction following, We assess our models using two of the most popular open-ended instruction-following benchmarks: AlpacaEval 2 (Li et al., 2023) and Arena-Hard v0.1 (Li et al., 2024). Both benchmarks evaluate the models’ versatile conversational abilities across a diverse set of queries. For each query, the evaluated model’s response and reference model’s response are compared head-to-head using an auto-evaluator. We use the officially recommended configurations111AlpacaEval: https://github.com/tatsu-lab/alpaca_eval. Arena-Hard v0.1: https://github.com/lm-sys/arena-hard-auto. during the evaluation.
Baselines.
We compare our approach with various powerful offline preference optimization methods, including FIGA (Guo et al., 2023), DPO (Rafailov et al., 2024b), and DPO variants (IPO (Azar et al., 2024), KTO (Ethayarajh et al., 2024), ORPO (Hong et al., 2024), TDPO (Zeng et al., 2024), R-DPO (Park et al., 2024), and SimPO (Meng et al., 2024)). The training objectives of these methods are listed in Table 6.
5 Experimental Results
In this section, we present the main results of our experiments, showcasing the superior performance of our method across various benchmarks and ablation studies (§5.1). Next, we conduct in-depth quantitative analyses to elucidate why our method outperforms DPO (§5.2 and §5.3). Furthermore, we demonstrate the versatility of our framework by adapting it to other DPO variants (§5.4).
5.1 Main results and ablations
Our method consistently and significantly outperforms baselines.
As shown in Table 1, our model DPO-BMC achieves the best overall performance across all benchmarks on QA and math tasks. Notably, DPO-BMC outperforms DPO by 3.8 absolute points on QA tasks and by 1.3 points on math tasks. On instruction-following tasks, as shown in Table 2, DPO-BMC secures the highest length-controlled win rate, surpassing DPO by over 5 points across various settings. The length-controlled win rate (Dubois et al., 2024) is a debiased version of the win rate, significantly reducing length gameability in LLM-based auto-annotation. It can be seen that while achieving a higher raw win rate, DPO-BMC’s average response length is notably shorter than that of other baselines (a case study is presented in Table 8). This attribute is credited to the correlated preference data we constructed, which directs optimization towards critical desired behaviors rather than verbosity. Although FIGA aims to steer the model in a similar direction, it fails to improve performance over the SFT model, as it focuses solely on different tokens and employs fixed token-level rewards. These consistent and significant improvements underscore the robustness and effectiveness of our method.
| Method | Question-Answering Tasks | Mathematical Reasoning Tasks | ||||||||
| ECQA | QASC | OBQA | StrategyQA | Avg. | GSM8k | MATH | MAWPS | TabMWP | Avg. | |
| SFT | 72.8 | 54.5 | 51.8 | 56.9 | 59.0 | 55.8 | 11.6 | 80.3 | 42.8 | 47.6 |
| FIGA | 70.3 | 52.5 | 51.7 | 48.6 | 55.8 | 54.1 | 9.8 | 75.5 | 39.0 | 44.6 |
| IPO | 71.5 | 58.9 | 53.6 | 58.4 | 60.6 | 57.2 | 12.1 | 82.2 | 42.5 | 48.5 |
| KTO | 71.6 | 56.5 | 52.8 | 56.0 | 59.2 | 56.4 | 12.0 | 80.1 | 42.5 | 47.8 |
| OPRO | 69.8 | 55.1 | 51.4 | 57.2 | 58.4 | 56.0 | 12.4 | 80.8 | 41.3 | 47.6 |
| TDPO | 74.4 | 58.2 | 54.8 | 59.1 | 61.6 | 57.3 | 12.2 | 81.6 | 43.2 | 48.6 |
| R-DPO | 73.5 | 59.5 | 55.4 | 58.8 | 61.8 | 56.9 | 12.0 | 81.9 | 42.2 | 48.2 |
| SimPO | 71.9 | 56.7 | 52.2 | 55.4 | 59.1 | 57.5 | 12.7 | 81.8 | 43.5 | 48.9 |
| DPO | 73.1 | 58.8 | 55.6 | 57.8 | 61.3 | 56.3 | 12.3 | 81.2 | 43.4 | 48.3 |
| DPO-BMC | 75.9 | 63.0 | 60.4 | 61.0 | 65.1 | 58.4 | 13.0 | 83.1 | 43.8 | 49.6 |
| DPO-BC | 75.7 | 62.0 | 56.0 | 60.1 | 63.4 | 57.6 | 12.7 | 82.8 | 43.4 | 49.1 |
| DPO-MC | 74.8 | 60.0 | 56.4 | 58.8 | 62.5 | 57.2 | 12.5 | 82.4 | 43.0 | 48.8 |
| Method | Llama3-Base (8B) | Mistral-Base (7B) | ||||||||
| AlpacaEval 2 | Arena-Hard | AlpacaEval 2 | Arena-Hard | |||||||
| LC (%) | WR (%) | Avg. len | WR (%) | Avg. len | LC (%) | WR (%) | Avg. len | WR (%) | Avg. len | |
| SFT | 7.5 | 4.7 | 956 | 2.6 | 414 | 8.1 | 5.9 | 998 | 2.2 | 454 |
| FIGA | 8.4 | 4.2 | 1,199 | 5.1 | 416 | 7.0 | 4.9 | 1,378 | 2.5 | 461 |
| IPO | 13.4 | 9.8 | 1,430 | 14.0 | 477 | 12.5 | 10.8 | 1,588 | 8.5 | 522 |
| KTO | 13.7 | 12.2 | 1,633 | 12.9 | 547 | 13.3 | 10.6 | 1,487 | 8.2 | 543 |
| ORPO | 12.5 | 11.4 | 1,793 | 11.7 | 573 | 14.5 | 11.5 | 1,630 | 9.4 | 566 |
| TDPO | 16.5 | 15.3 | 1,766 | 16.4 | 564 | 14.8 | 12.8 | 1,740 | 11.2 | 551 |
| R-DPO | 17.1 | 14.4 | 1,801 | 17.6 | 582 | 16.0 | 12.3 | 1,521 | 10.4 | 529 |
| SimPO | 21.3 | 18.9 | 1,718 | 26.6 | 562 | 16.8 | 14.4 | 1,906 | 18.4 | 615 |
| DPO | 16.0 | 14.8 | 1,713 | 17.6 | 559 | 15.1 | 13.3 | 1,657 | 13.6 | 540 |
| DPO-BMC | 22.4 | 16.8 | 1,285 | 18.1 | 406 | 20.8 | 16.6 | 1,317 | 17.6 | 488 |
| DPO-BC | 20.6 | 14.4 | 1,269 | 16.8 | 422 | 18.6 | 13.8 | 1,489 | 15.9 | 502 |
| DPO-MC | 17.7 | 15.2 | 1,890 | 17.9 | 579 | 16.4 | 14.3 | 1,712 | 15.4 | 551 |
| Method | QA | Math | IF |
| SFT | 56.9 | 47.6 | 7.5 |
| IPO | 60.6 | 48.3 | 13.4 |
| IPO-BMC | 64.1 | 48.6 | 15.7 |
| ORPO | 58.4 | 47.6 | 12.5 |
| ORPO-BMC | 62.3 | 48.4 | 15.7 |
| R-DPO | 61.8 | 48.2 | 17.1 |
| R-DPO-BMC | 65.3 | 49.5 | 20.0 |
| SimPO | 59.1 | 48.9 | 21.3 |
| SimPO-BMC | 61.6 | 49.0 | 19.7 |
| DPO | 61.3 | 48.3 | 16.0 |
| DPO-BMC | 65.1 | 49.6 | 22.4 |
Both key designs in BMC are crucial.
In Table 1 and Table 2, we additionally present results from ablating each key design element of DPO-BMC:
DPO-BC: Training using DPO’s original objective on our constructed preference data.
DPO-MC: Training using our proposed objective in Eq. 5 on the original preference data.
Our examination reveals several key findings: (1) enhancing the correlation between preference pairs remarkably boosts model performance; (2) even when using the same training preference data, our designed optimization objective outperforms DPO, highlighting its superior ability to model fine-grained correlations; and (3) combining our constructed data with our designed objective yields the best results, affirming the inseparability of the Bridging Phase and the Modeling Phase—bridging correlation and then modeling it.
Influence of .
Here we investigate the impact of the threshold in the DPO-BMC objective on model performance, as illustrated in Figure 3. Setting reduces our method to one that assigns fixed token-level rewards. The optimal value of varies across different tasks, highlighting the need for task-specific tuning. Additionally, a large value can degrade model performance due to excessively aggressive gradient updates on certain tokens, which negatively affects both QA and instruction-following tasks.
5.2 Quantitative analysis of bridging and modeling phase
To rigorously assess the effectiveness of the two pivotal phases in our framework, we segment the 60k training data of UltraFeedback into six equal-sized splits, ordered by increasing edit distance between the winning and losing responses. For each split, we also construct its corresponding pair data through our Bridging Phase. We then train four models—–(a) DPO, (b) DPO-MC, (c) DPO-BC, and (d) DPO-BMC—on each split, and report the length-controlled win rate on AlpacaEval 2 and the average gradient norm during training. All models are trained with identical hyperparameters to ensure consistency and comparability. It can be observed from Figure 4 that the Bridging Phase successfully decreases the edit distance between pairwise data through targeted modification, shifting the optimization process to concentrate on the most critical differences in preference data. The results underscore the efficacy of the Bridging Phase in consistently enhancing the performance across all splits, signifying its role in refining the training data for more focused learning. A notable observation is the behavior of the average gradient norm during the training of DPO, which exhibits an increasing trend as the edit distance between pairwise data enlarges. This trend highlights the sensitivity of DPO’s training process to individual data points, which can potentially lead to gradient imbalances and inefficient learning. However, the proposed learning objective in our Modeling Phase effectively mitigates these imbalances through a dynamic training process that adjusts according to the policy model’s confidence. This adaptive mechanism not only ensures that the model places greater emphasis on learning challenging distinctions but also reinforces the patterns that have already been acquired, culminating in a more nuanced and robust policy. Such an approach is crucial in achieving a balanced optimization landscape with diverse training data, leading to superior generalization and performance in downstream tasks.
In Figure 5, we present the KL divergence between the policy model trained with DPO, DPO-MC, DPO-BC, and DPO-BMC with identical hyperparameters and the reference model, measured on the winning responses from a held-out set of UltraFeedback during training. The results also validate our above analyses: (1) the Bridging Phase fosters tailored learning toward critical differences in preference data, potentially resulting in a larger KL divergence during training; (2) our meticulously designed loss function in the Modeling Phase effectively moderates the optimization intensity across diverse training data, thereby achieving a more controlled and steady KL divergence.
5.3 Quantitative analysis of credit assignment
We quantitatively compare the token-level and sequence-level credits assigned by DPO and DPO-BMC, assessing how well their final learned rewards align with preference labels on a held-out set of UltraFeedback.
Analysis on token-level reward.
Figure 6 depicts the token-level reward assignment for DPO and DPO-BMC on a response pair consisting of a preferred response and a dispreferred response . The reward of each token is computed as . From the figure, we observe that: (1) DPO assigns nearly uniform rewards to each token within a response, failing to differentiate the importance of tokens to the overall response quality; and (2) while DPO can identify and assign lower rewards to several erroneous tokens in the dispreferred response, such as “13” and assigning them small rewards, it struggles to discriminate subtle differences between the preferred and dispreferred responses. In contrast, DPO-BMC not only assigns relatively higher rewards to critical tokens like “descending order” but also precisely identifies the incorrect tokens inside the dispreferred response. Notably, within the incorrect span of “22, 13, 99”, DPO-BMC assigns a significantly low reward to the beginning token “22” while imposing lesser penalties on the subsequent ones. This pattern of reward assignment is more reasonable as it reflects the autoregressive nature of the generation, where subsequent mistakes are conditioned on previous errors. This example highlights the limitation of DPO in providing precise token-level preferences on sentence quality, and our method can effectively alleviate this issue. As illustrated, a generated sequence may contain a mix of correct and incorrect information, or even accurate and helpful responses marred by minor errors. To this end, our method provides a more effective and interpretable approach to discerning the subtle quality of these outputs.
Analysis on sequence-level reward.
For a rigorous comparison, we calculate the sequence-level DPO reward expression by (with the partition function excluded) and present the results in Table 4. The reward margin is determined by . Reward accuracy is defined as the percentage of preference pairs where the winning response achieves a higher reward than the losing response, i.e., . Our findings demonstrate that DPO-BMC outperforms DPO in terms of average reward margin and reward accuracy. This enhancement validates the superior ability of our method to discern subtle differences between preference pairs, thereby facilitating more effective generalization.
| Method | Avg. Reward Margin | Reward Accuracy (%) |
| DPO | 0.54 | 72.19 |
| DPO-BMC | 0.74 | 73.60 |
5.4 Versatility of our framework
Our BMC framework is versatile and can be applied to various DPO variants. Table 3 shows the performance comparison between the original XPO and XPO-BMC methods. Given that KTO only requires a binary signal to indicate whether an output is desirable or undesirable for a given input, and TDPO is based on token-level MDP, we exclude these from our consideration. The results indicate that XPO-BMC variants generally surpass their corresponding XPO methods across various tasks. Specifically, IPO-BMC, ORPO-BMC, R-DPO-BMC, and DPO-BMC achieve higher average accuracies in QA and Math, as well as superior LC scores on AlpacaEval 2 for IF. However, SimPO-BMC does not outperform SimPO on the IF task, showing a slight decrease from 21.3 to 19.7. These findings underscore the robustness and efficacy of our BMC approach in enhancing the performance of different DPO methods, affirming its potential for broad application.
6 Conclusion
In this work, we propose BMC, an effective framework for bridging and modeling correlations in pairwise data for direct preference optimization. BMC equips LLMs with better human value alignment through a two-phase process: a Bridging Phase that enhances correlations between pairwise data by explicitly manifesting fine-grained preference signals via targeted modifications, and a Modeling Phase that learns token-level correlations by dynamically leveraging the the policy model’s confidence during training. Our framework exhibits superior performance in question-answering, mathematical reasoning, and instruction-following tasks, consistently surpassing the baseline DPO by a significant margin. Extensive analysis highlights that the key designs in BMC are crucial and validates the effectiveness and versatility of BMC.
Appendix A Appendix
A.1 Data statistics and evaluation metrics used for experiments
We list the detailed data statistics and evaluation metrics of our experiments in Table 5. Our experiments comprise both closed-ended evaluation (QA and math) and open-ended evaluation (instruction following).
| Task | Train / Test | Dataset | Number | Metric |
| QA | Train | ECQA | 7,598 | |
| QASC | 8,134 | |||
| Test | ECQA | 2,194 | Accuracy | |
| QASC | 926 | |||
| OBQA | 500 | |||
| StrategyQA | 687 | |||
| Math | Train | MetaMathQA | 40,000 | |
| Test | GSM8k | 1,319 | Accuracy | |
| MATH | 5,000 | |||
| MAWPS | 2,065 | |||
| TabMWP | 1,000 | |||
| IF | Train | UltraFeedback | 61,135 | |
| Test | AlpacaEval 2 | 805 | Win rate | |
| Arena-Hard | 500 |
A.2 Prompt template for targeted modification
We demonstrate the prompt template of targeted modification for question answering and mathematical reasoning tasks in Figure 7. Since the SFT model has been fine-tuned on the ground truth for QA and math tasks, the inferred output may be quite approximate to in some circumstances. Therefore, we require the off-the-shelf LLM to filter out preference pairs where is good enough. Finally, we filtered out 31% data and 43% data for the QA task and math task, respectively. Note that for the training data of our baselines like DPO, we also use the filtered pairs for a fair comparison. For instruction-following tasks, the prompt template we use is shown in Figure 8.
A.3 Implement details
To ensure a fair comparison, we conduct thorough hyperparameter tuning for all methods compared in our experiments.
SFT training hyperparameters.
We train SFT models using the following hyperparameters: a learning rate of 2e-5, a batch size of 128, a max sequence length of 2048, and a cosine learning rate schedule with 10% warmup steps. For QA and instruction-following tasks, we train the model for 1 epoch, whereas for mathematical tasks, we extend the training to 2 epochs. All the models are trained with an Adam optimizer (Kingma & Ba, 2014).
Preference optimization training hyperparameters.
During preference optimization, we performed initial experiments to determine the optimal batch sizes in [32, 64, 128] and training epochs in [1, 2, 3]. Our results indicate that using a batch size of 128 and a single training epoch consistently produces the best outcomes across all methods. Consequently, we adopted these parameters for all subsequent preference optimization experiments. We also configured the maximum sequence length to 2048 and employed a cosine learning rate schedule with a 10% warmup period for training on the preference optimization dataset. For method-specific training hyperparameters, we individually search the learning rates in the range of [3e-7, 5e-7, 6e-7, 1e-6] for each method. Besides, we conduct a grid search according to Table 6 and report the best performance. Table 7 shows the hyperparameters of our method used under each setting.
| Method | Objective | Hyperparameter |
| FIGA | ||
| IPO | ||
| KTO | ||
| where | ||
| ORPO | ||
| where | ||
| TDPO | ||
| R-DPO | ||
| SimPO | ||
| DPO | ||
| DPO-BMC | ||
| where | ||
| Task | Model | Learning Rate | ||
| QA | Llama2-7B-base | 5e-7 | 0.05 | 3.0 |
| Math | Llama2-7B-base | 5e-7 | 0.05 | 2.5 |
| IF | Llama3-8B-base | 5e-7 | 0.01 | 1.5 |
| Mistral-7B-base | 5e-7 | 0.01 | 2.0 |
A.4 Case Study
In this case study, we analyze the performance of various methods, all based on the Llama3-8B-base model, using a sample instruction from AlpacaEval 2. Table 8 displays the responses generated by four different methods. The SFT model produces a concise and factually accurate answer, although it lacks detail and context regarding the name change. In contrast, models further trained with SimPO and DPO yield incorrect responses, asserting that Facebook Corporation did not change its legal name and providing an inaccurate account of the rebranding process. Our proposed method, DPO-BMC, demonstrates a notable improvement by providing a response that is both factually accurate and comprehensive. Additionally, the response generated by our method is significantly shorter than those produced by DPO and SimPO. This combination of detail, accuracy, and brevity highlights the superiority of our approach in delivering informative and precise answers.
| Instruction | Did Facebook corporation change its name? |
| SFT | Yes, Facebook Corporation changed its name to Meta Platforms, Inc. In 2021. |
| SimPO | No, the Facebook Corporation did not alter its legal name. Despite the fact that the company has introduced a diverse array of products and services beyond its primary social networking platform, which includes popular applications such as Instagram, WhatsApp, Messenger, and various others, the official name of the corporation continues to be Facebook, Inc. Nevertheless, in October 2021, Facebook made an announcement indicating their intention to modify their corporate brand name to better reflect their expanding range of offerings. As of now, the new brand name has not been disclosed to the public, and it is anticipated that the complete transition to this new name will require a significant amount of time. |
| DPO | No, Facebook Corporation did not change its name. Facebook Corporation, also known as Facebook, Inc., is the parent company of the Facebook social media platform, and it has not gone through any significant name change. The company holds other online platforms and services such as Instagram and WhatsApp, but the name Facebook Corporation remains the same. There have been some changes in the branding, such as the rebranding of WhatsApp and Facebook Messenger with the Meta logo, but the parent company’s name has not changed. The announcement made by Facebook in October 2021 about rebranding its parent company as Meta did not materialize, and Facebook Corporation still holds the name Facebook Corporation. |
| DPO-BMC | Yes, Facebook Corporation changed its name to Meta Platforms, Inc. In October 2021. The rebranding aims to reflect the company’s expanded focus beyond social media and into building the ”metaverse,” a virtual space where people can interact and communicate in immersive, 3D environments. The new name ”Meta” represents the company’s ambition to drive breakthroughs in areas such as virtual and augmented reality. Despite the name change, Facebook remains the name of the company’s flagship social network. |